OpenStudy (anonymous):

how many 1\4 are in 5\6

OpenStudy (phi):

divide 1/4 into 5/6 \[ \frac{\frac{5}{6}}{\frac{1}{4}} \] to simpliy that multiply top and bottom by 4 \[ \frac{\frac{5}{6}}{\frac{1}{4}} \cdot \frac{4}{4} = \frac{\frac{5}{6}\cdot 4}{\frac{1}{4}\cdot 4}\]

OpenStudy (phi):

the bottom of the complicated fraction is 1/4 *4 what is 1/4 * 4 ?